Please show all work!!!
11.* A random sample of size n 64 is drawn from a population with mean μ and standard deviation σ. The mean and standard deviation of the sample are X = 308.9 and s 31.9 a. Find a 90%confidence interval for the mean μ. Interpret this interval. b. Find a 95%confidence interval for the mean μ. Interpret this interval. c. Find a 99%confidence interval for the mean μ. Interpret this interval. d. Compare the widths of the three confidence intervals. Which interval is the widest? The narrowest?
Answer #1

This is the case of finding the confidence interval for the mean \mu when \sigma^2 is unknown,

and this confidence interval is given by,

P[\bar{x}-\frac{s}{\sqrt{n}}t_{\alpha/2}\leq \mu \leq \bar{x}+\frac{s}{\sqrt{n}}t_{\alpha/2}]=1-\alpha

where,

\bar{x}=sample\quad mean

s=sample\quad standard \quad deviation

n=total\quad observations

we are given that,

\bar{x}=308.9

s=31.9 and

n=64

therefore 100(1-\alpha)% % is given by,

P[308.9-\frac{31.9}{\sqrt{64}}t_{\alpha/2}\leq \mu \leq 308.9+\frac{31.9}{\sqrt{64}}t_{\alpha/2}]=1-\alpha

(a).

for a 90% confidence interval \alpha=0.1

\therefore t_{\alpha/2}=t_{0.05}=1.671[value generated from the t-table]

hence the required confidence interval is given by,

[308.9-\frac{31.9}{8}*1.671,308.9+\frac{31.9}{8}*1.671]

=[308.9-6.663,308.9+6.663]

=[302.237,315.563]

That is we can say with 90% confidence that the true value of the mean lies between [302.237,315.563]

(b).

for a 95% confidence interval \alpha=0.05

\therefore t_{\alpha/2}=t_{0.025}=2[value generated from the t-table]

hence the required confidence interval is given by,

[308.9-\frac{31.9}{8}*2,308.9+\frac{31.9}{8}*2]

=[308.9-7.975,308.9+7.975]

=[300.925,316.875]

That is we can say with 95% confidence that the true value of the mean lies between [300.925,316.875]

(c).

for a 99% confidence interval \alpha=0.01

\therefore t_{\alpha/2}=t_{0.005}=2.660[value generated from the t-table]

hence the required confidence interval is given by,

[308.9-\frac{31.9}{8}*2.66,308.9+\frac{31.9}{8}*2.66]

=[308.9-10.61,308.9+10.61]

=[298.29,319.51]

That is we can say with 99% confidence that the true value of the mean lies between [298.29,319.51]

(d).

Now comparing the three confidence interval

a - [302.237,315.563]

b - [300.925,316.875]

c - [298.29,319.51]

the widest interval is the 99% confidence interval that is (c) - [298.29,319.51]

and the narrowest interval is the 90% confidence interval that is (a) - [302.237,315.563]